OpenStudy (anonymous):

help please: If f(x,y) = x²+xy²-y³ and x=t²+3 and y=t³+5t , then find δf/δx δf/δt

OpenStudy (anonymous):

\[\frac{∂f}{∂x}=\frac{∂}{∂x}(x^2+xy^2-y^2)\]Just treat the y's retriceconstants. For the second one you need to use the chain rule for functions of multiple variables. \[\frac{∂f}{∂t}=\frac{∂f}{∂x} \frac{∂x}{∂t}+\frac{∂f}{∂y}\frac{∂y}{∂t}=f_xx_t+f_yy_t\]

OpenStudy (anonymous):